# Efficient Implementation of Gaussian Process Regression Accelerated Saddle Point Searches with Application to Molecular Reactions

### v1

@@ 3.1 Savings by using GPR
Figure 1 shows a comparison of the number of HF calculations needed to reach convergence with and without the use a PR. The same MMF method with the dimer for finding the minimum mode is used in both cases, but for the GPR-dimer, these calculations are carried out on the surrogate energy surface. An order of magnitude reduction in the HF calculations is obtained by using the GPR. The median number of calculations to convergence drops from 308 to 31. While HF calculations are relatively fast, this difference in the number of electronic structure calculations translates to large savings in computational effort when higher-level methods are used. The GPR surrogate energy surface construction and overhead is independent of the underlying method for generating the data, and so the computational cost is the same as long as atomic forces can be obtained.
[…]
@@
Figure 1: (a) Comparison of the median number of HF electronic structure calculations needed to converge on a saddle point, with and without the GPR acceleration, in minimum mode following based on the dimer method applied to the 500 system test set of hermesSellaOpenSourceAutomationFriendly2022 [18]. The median number of HF calculations drops from 308 to 31 by using the GPR, an order of magnitude reduction. (b) Number of systems where the difference between the energy of the initial atomic structure and that of the converged saddle point is within each of the 2 eV intervals shown on the horizontal axis. In almost all cases, the saddle point has lower energy than the initial configuration, most often within 2 eV, but there are some examples where the difference is even larger than 10 eV.
@@
More detailed information on this is shown in Figure 2. There the focus is on systems where the energy difference between saddle points obtained with and without the use of the GPR is less than 0.01 eV, a total of 373 cases. The systems are arranged according to the minimized RMSD of the Cartesian coordinates in the initial and converged configurations of the atoms. The range is from 0 to 30 Å. The savings in the number of HF calculations by using the GPR is clearly greater for larger RMSD values. This is to be expected, since the reduction in the number of HF calculations comes from being able to take several steps on the GPR surface without having to resample from the true energy surface. The GPR-dimer algorithm effectively dynamically adjusts the step size as seen from the prespective of the sampling of the true energy surface. The median number of HF calculations for this set is 29 when using the GPR-dimer while it is 301 withou the use of the GPR. The average reduction is 271 HF calculations. However, the distribution is broad, for three cases, where the RMSD is between 10 and 30 Å, the reduction in the number of calculations is over 1500. The savings by using a GPR is particularly large when the initial guess of the saddle point is poor.
@@
Figure 2: (a) Difference between the number of HF electronic structure calculations needed to converge on a saddle point with and without the GPR acceleration for each of the systems in the dataset, ordered along the horizontal axis according to the root mean square distance between the initial configuration of the atoms and that of the saddle point found. With the GPR acceleration the median is 29 compared to 301 without it. The mean reduction is indicated by the blue-gray line, 271 HF calculations. Calculations where the two methods give the same saddle point energy to within 0.01 eV are included, a total of 373 systems. (b) The RMSD distance (in Å) between the initial configuration of the atoms and that of the converged saddle point. Comparison with (a) shows how the efficiency of the GPR acceleration increases with the RMSD distance.
@@ 3.2 Comparing GPRD and Sella
Figure 3: Comparison of the number of HF electronic structure calculations needed to reach convergence using the GPR-dimer method which uses Cartesian coordinates and GPR acceleration, and the Sella method which makes use of internal coordinates. The number of HF calculations is on average similar for both methods, with the GPR-dimer requiring a median of 29 HF calculations while Sella requires 31. The comparison includes 345 systems where the two methods lead to the same saddle point energy to within 0.01 eV. In several cases, the two methods do not converge to the same saddle point, as illustrated in Figure 4.
There are several casess where the MMF calculation, with or without GPR, does not converge on the same saddle point as the calcuation with Sella. This happens, for example, for the first system in the dataset, singlet 000, as illustrated in Figure 4. The system in question is a 16-atom molecule with a stoichiometry of $\mathrm{C}_{5}\mathrm{OH}_{1}0$. The initial configuration of the atoms corresponds to an acyclic ether with a fairly large separation between the carbon endpoints of around 2.2 Å. As can be seen in Figure 4, the saddle point found by the GPR-dimer corresponds to the hydrogen from one of the endpoints transitioning to the other. At the saddle point, the C-H bond distances are $\approx 1.25$ and the nergy has dropped by 0.9 eV. The RMSD between the initial configuration of the atoms and the saddle point is only 0.2 Å. The Sella calculation comes close to this configuration during the course of the saddle point search, but instaed of converging, it proceeds to the final state of the transition and eventually converges on a saddle point for an irrelevant rotation of a methyl group, 0.4 eV lower in energy than the sadlde point found by the GPR-dimer method. The RMSD between this saddle point and the initial configuration is only 0.6 Å. The GPR-dimer calculation requires 23 HF calculations while the Sella calculation requires 197. This example illustrates how important it is to check the configuration of atoms obtained in a saddle point search. An NEB calculation of the minimum energy path connecting the initial configuration and the saddle point obtained with Sella shows that the saddle point obtained with the GPR-dimer is indeed located along the path and that the final state of the corresponding transition is an intermediate minimum before the metly group starts rotating in the Sella caluclation (see inset in Figure 4).
@@
The use of internal coordinates is particularly challenging when a geometry near a linear arrangement of three or more atoms is encountered. There, ghost atoms are introduced algorithmically in Sella to avoid singlularities in the internal coordinates. For the 8 systems in the dataset in Figure 2 where this occurs, the GPR-dimer requires 32 HF calculations, close to the average, while Sella requires 44, which is 42% larger than its average.
The first system within the dataset that involves the introduction of ghost atoms is system singlet 016. It involves two fragments, a methyl group ($\mathrm{CH}_{3}$) and a molecule with two ether linkages ($\mathrm{C}_{4}\mathrm{H}_{7}\mathrm{O}_{2}$), i.e. 1-methoxy-2-(2-methoxyethoxy)ethane. The calculation with Sella first increases the distance between the two fragments, but then eventually attaches the methyl group to form a single chain. The saddle point search requires 114 HF calculations and the saddle point reached is 2.4 eV below the initial configuration with an RMSD of 4.2 Å. The GPR-dimer, however, converges after 42 HF calculations to a different saddle point where an H atom jumps onto the methyl group to form $\mathrm{C}_{4}$ and a $\mathrm{CO}_{2}$ dissociates from the molecule. This saddle point is 0.4 eV higher than the one found by Sella and the RMSD with respect to the initial configuration is 1.7 Å. As in the singlet 000 casse, the path taken by the Sella calculation comes close to the saddle point found by the GPR-dimer but proceeds to a saddle that is lower in energy and farther away from the initial guess for the saddle point configuration.
@@ 4 Discussion
The 500 system data set is constructed in such a way that it only includes cases where Sella converges to some saddle point [18]. For 10 of these systems, the GPR-dimer calculation ends up being aborted using one of the 3 criteria listed in Section 2. In most cases, this occurs because the electronic structure calculation fails and the NWChem software reports an error when atoms have come too close together during the saddle point search on the GPR surrogate energy surface. Also, several of the systems in the dataset involve two separate molecular fragments and in a few cases, the saddle point calculation moves them away from each other until the calculations are aborted. The reason for this is that the direction of the minimum mode then corresponds to the uniform displacement of the two fragments with little change in their internal structure. The energy decreases as the fragments move closer together and so the saddle point search moves the fragments in the opposite direction further apart so as to climb uphill on the energy surface. Similar behaviour is observed in the Sella calculations of these systems. However, in most cases where this occurs, the sideways minimization of the energy eventually leads to the search path turning around and the two fragments then eventually coming together in a chemical reaction. However, in a few cases, this does not happend and the GPR-dimer calculation is eventually aborted after a maximum number of iterations is reached. When the initial configuration of the atoms consists of separated molecular fragments, it could make more sense to start with a local minimization of the energy to bring the fragments to an optimal (relative) position, and then start the saddle point searches from configurations of the atoms that are oly slightly perturbed from the minima. Here, we have chosen to work with the input as presented in the literature.
@@
The application of saddle point search methods designed to move uphill in energy is actually questionable when the initial state consists of two molecular fragments since the initial state may not be well described with a harmonic approximation of the energy surface and harmonic transition state theory is not expected to provide a good estimate of the rate. An NEB calculation of the minimum energy path for such a reaction does not run into this kind of problem, but then the final state of the reaction needs to be specified. In principle, a free energy calculation should be applied in such cases, and, an optimal dividing surface should be determined by maximizing the free energy [21].
@@
An important aspect of the use of a surrogate energy surface is not to venture too far from the region of the data points in the training set since extrapolation is not reliable. This is implemented by early stopping criteria in the GPR-dimer saddle point searches [26, 25]. The number of times early stopping criteria correlates with longer searches, as the structures added to the training data due to the early stopping are not typically near a saddle point. Here, Euclidean distances have been used to decide on early stopping, but we expect that a Hausdorff distance measure could work better in tandem, and this will be tested in future refinements of the method. Additionally, the Euclidean distance measure is not sensitive to changes across chemical species, which can lead to chemically unfeasible structures being marked as within the cutoff.
@@
The GPR-dimer saddle point on the surrogate surface can lead to configurations of the atoms that have high energy. The NWChem software used for the HF calculations performs checks on the input structures before calculating the energy and forces, and these are triggered even when the GPR is not used. While the GPR-dimer explores unphysical structures, high energy and atomic forces are in principle not problematic; the GPR can successfully handle such regions. However, NWChem often terminates when atoms come too close and this prevents the GPR-dimer from navigating these regions. We hypothesise that by providing some energy/force value, even a large one representing extreme repulsion, as a pseudo-output from the electronic structure calculation through a wrapper in ASE or EON would allow the GPR-dimer to successfully converge to a saddle point in such cases, while here they represent a failure.
@@
Another improvement that is being pursued for future refinements of the method is the pruning of input data, namely, the elimination of some of the configurations that have been calculated with the electronic structure method early in the search and are far from the saddle point. This will help reduce the computational cost of the matrix inversion order and the memory requirement.
@@
In future work we intend to apply these and other methods to larger and more diverse benchmark sets that are not biased towards any particular methods. The results of such studies will be reported in an ongoing collaboration to extend the OptBench suite of benchmarks for finding saddle points and minimum energy paths of thermally activated transitions [4].
@@ 5 Conclusions
The results presented here demonstrate that saddle point searches for molecular reactions can be performed efficiently with Cartesian coordinates despite the large disparity in the stiffness of the molecular degrees of freedom if the calculation makes use of GPR to generate a surrogate energy surface. The MMF method is used here in combination with a dimer estimate of the minimum mode. The use of a GPR reduces the number of electornic structure calcuations needed to reach convergence by an order of magnitude.
The calculations are carried out for a dataset of 500 systems presented previously by [18]. It consists of guesses for saddle points and by construction only includes cases where calculations using Sella are successful in in that some saddle point is obtained. A comparison is made between the performance of the GPR-dimer method using Cartesian coordinates and the algorithm in Sella, which is based on internal coordinates and the performance of the two found to be comparable in terms of the HF calculations. However, in several cases the Sella calculations bypass the saddle point closest to the initial guess and converges instead on some rearrangement of the atoms in the final state, such as a rotation of a methyl group. This demonstrates the importance of checking the saddle point configuration of atoms obtained in a single endpoint calculation. This can be done efficiently by using the NEB method which reveals if the saddle point found is adjacent to the initial state configuration or if there are other saddle points present in between.
While the use of a GPR presents a significant overhead for the MMF calculation, the efficient implementation presented here in C++ leads to a reduction in the overall computational time even though the electronic structure calculations are carried out here at a low HF level. When higher level electronic structure calculations are used, the overhead will be an even smaller fraction of the total computational effort. The scaling of the computational effort in the generation of the GPR surrogate energy surfaces with the dimensionality of the system and the number of electronic structure results included in the training set is, however, rapid and further optimisation is required in order to be able to apply this approach to large systems. Pruning of the input data to drop points far from the saddle point region is one option and the use of only a subset of the atom coordinates of a large system focusing thereby on the region of interest is another, possibly followed by later refinement of the saddle point found with respect to all coordinates.
Acknowledgments
@@
From a numerical standpoint, all code and data for reproducing the analysis are provided in a single repository, which is tagged. These include pinned versions of ase, EON (from https://theochemui.github.io/eOn/), the GPR dimer code, and a copy of the spack recipes [10] repository for building NWChem used for the Hartree-Fock calculations along with spin-out utilities from Wailord [12]. A conda ecosystem-based environment managed by pixi is provided to further facilitate reproducibility, with the modules either pinned to versions or bundled within the same repository via git-subrepo following best practices for literate programming

### v2

@@ 3.1 Savings by using GPR
Figure 1 shows a comparison of the number of HF calculations needed to reach convergence with and without the use of GPR. The same MMF method with the dimer for finding the minimum mode is used in both cases, but for the GPR-dimer, these calculations are carried out on the surrogate energy surface. An order of magnitude reduction in the number of HF calculations is obtained by using the GPR. The median number of calculations needed for convergence drops from 308 to 31. While HF calculations are relatively fast, this difference in the number of electronic structure calculations translates to large savings in computational effort when higher-level methods are used. The GPR surrogate energy surface construction and overhead are independent of the electronic structure method used for generating the data.
In general, the initial guesses of saddle point configurations of the atoms in the Sella benchmark have higher energy than the saddle points. This is unusual as saddle point searches are more often started near an energy minimum corresponding to an initial state and the saddle point then approached from below. The results in Figure 1 are arranged in a histogram according to the energy difference between the initial configuration and the obtained saddle point. This energy difference can even be larger than 10 eV for a couple of systems, but these are excluded from the graphic for visual clarity. The number of systems in the data set binned in 2 eV intervals is shown in Figure 1 (b). Most systems start within 2 eV above the saddle point. Only a few systems start from a configuration that has lower energy than the saddle point. The two sets of calculations, with and without the use of a GPR, most often converge on the same saddle point, but not always, as can be discerned from the color coding of Figure 1 (b). For example, there are more systems that end up on a saddle point that is within 2 eV of the starting point when the GPR is not used, while there are more systems that end up between 4 and 6 eV below the initial point in the GPR-dimer calculations.
@@
Figure 1: (a) Comparison of the median number of HF electronic structure calculations needed to converge on a saddle point, with and without the GPR acceleration, in minimum mode following calculations based on the dimer method applied to the 500 system test set of [23]. The median number of HF calculations drops from 308 to 31 by using the GPR. (b) Number of systems where the difference between the energy of the initial atomic structure and that of the converged saddle point is within each of the 2 eV intervals shown on the horizontal axis. In almost all cases, the saddle point has lower energy than the initial configuration, most often within 2 eV, but there are some examples where the difference is even larger than 10 eV.
@@
More detailed information on this is shown in Figure 2. There the focus is on systems where the energy difference between saddle points obtained with and without the use of the GPR is less than 0.01 eV, a total of 373 cases. The systems are arranged according to the minimized RMSD of the Cartesian coordinates in the initial and converged configurations of the atoms. The range is from 0 to 30 Å. The savings in the number of HF calculations by using the GPR is clearly greater for larger RMSD values. This is to be expected, since the reduction in the number of HF calculations comes from being able to take several steps on the GPR surface without having to resample from the true energy surface. The GPR-dimer algorithm effectively dynamically adjusts the step size as seen from the perspective of the sampling of the true energy surface. The median number of HF calculations for this set is 29 when using the GPR-dimer while it is 301 without the use of the GPR. The average reduction is 271 HF calculations. However, the distribution is broad, for three cases, where the RMSD is between 10 and 30 Å, the reduction in the number of calculations is over 1500. The savings by using a GPR is particularly large when the initial guess of the saddle point is poor.
@@
Figure 2: (a) Difference between the number of HF electronic structure calculations needed to converge on a saddle point with and without the GPR acceleration for each of the systems in the dataset, ordered along the horizontal axis according to the root mean square distance between the initial configuration of the atoms and that of the saddle point found. With the GPR acceleration the median is 29 compared to 301 without it. The mean reduction is indicated by the blue-gray line, 271 HF calculations. Only calculations where the two methods give the same saddle point energy to within 0.01 eV are included, a total of 373 systems. (b) The RMSD distance (in Å) between the initial configuration of the atoms and that of the converged saddle point. Comparison with (a) shows how the efficiency of the GPR acceleration increases with the RMSD distance.
@@ 3.2 Comparing GPRD and Sella
Figure 3: Comparison of the number of HF electronic structure calculations needed to reach convergence using the GPR-dimer method which uses Cartesian coordinates in combination with GPR acceleration, and the Sella method which makes use of internal coordinates. The number of HF calculations is on average similar for both methods, with the GPR-dimer requiring a median of 29 HF calculations while Sella requires 31. The comparison includes 345 systems where the two methods lead to the same saddle point energy to within 0.01 eV. In several cases, the two methods do not converge to the same saddle point, as illustrated in Figure 4.
There are several cases where the MMF calculation, with or without GPR, does not converge on the same saddle point as the calculation with Sella. This happens, for example, for the first system in the dataset, singlet 000, as illustrated in Figure 4. The system in question is a 16-atom molecule with a stoichiometry of $\mathrm{C}_{5}\mathrm{OH}_{10}$. The initial configuration of the atoms corresponds to an acyclic ether with a fairly large separation between the carbon endpoints of around 2.2 Å. As can be seen in Figure 4, the saddle point found by the GPR-dimer corresponds to the hydrogen from one of the endpoints transitioning to the other. At the saddle point, the C-H bond distances are $\approx 1.25$ and the energy has dropped by 0.9 eV. The RMSD between the initial configuration of the atoms and the saddle point is only 0.2 Å. The Sella calculation comes close to this configuration during the course of the saddle point search, but instead of converging, it proceeds to the final state of the transition and eventually converges on a saddle point for an irrelevant rotation of a methyl group, 0.4 eV lower in energy than the saddle point found by the GPR-dimer method. The RMSD between this saddle point and the initial configuration is 0.6 Å. The GPR-dimer calculation requires 23 HF calculations while the Sella calculation requires 197. This example illustrates how important it is to check the configuration of atoms obtained in a saddle point search. An NEB calculation of the minimum energy path connecting the initial configuration and the saddle point obtained with Sella shows that the saddle point obtained with the GPR-dimer is indeed located along the path and that the final state of the corresponding transition is an intermediate minimum before the methyl group starts rotating in the Sella calculation (see inset in Figure 4).
@@
The use of internal coordinates is particularly challenging when a geometry near a linear arrangement of three or more atoms is encountered. There, ghost atoms are introduced algorithmically in Sella to avoid singularities in the internal coordinates. For the 8 systems in the dataset in Figure 2 where this occurs, the GPR-dimer requires 32 HF calculations, close to the average, while Sella requires 44, which is 42% larger than its average.
The first system within the dataset that involves the introduction of ghost atoms is system singlet 016. It involves two fragments, a methyl group ($\mathrm{CH}_{3}$) and a molecule with two ether linkages ($\mathrm{C}_{4}\mathrm{H}_{7}\mathrm{O}_{2}$), i.e. 1-methoxy-2-(2-methoxyethoxy)ethane. The calculation with Sella first increases the distance between the two fragments, but then eventually attaches the methyl group to form a single chain. The saddle point search requires 114 HF calculations and the saddle point reached is 2.4 eV below the initial configuration with an RMSD of 4.2 Å. The GPR-dimer, however, converges after 42 HF calculations to a different saddle point where an H atom jumps onto the methyl group to form %$\mathrm{C}_{4}$ $\mathrm{CH}_{4}$ and a $\mathrm{CO}_{2}$ dissociates from the molecule. This saddle point is 0.4 eV higher than the one found by Sella and the RMSD with respect to the initial configuration is 1.7 Å. As in the singlet 000 case, the path taken by the Sella calculation comes close to the saddle point found by the GPR-dimer but proceeds to a saddle that is lower in energy and farther away from the initial guess for the saddle point configuration.
@@ 4 Discussion
The 500 system data set is constructed in such a way that it only includes cases where Sella converges to some saddle point [23]. For 10 of these systems, the GPR-dimer calculation ends up being aborted using one of the 3 criteria listed in Section 2. In most cases, this occurs because the electronic structure calculation fails and the NWChem software reports an error when atoms have come too close together during the saddle point search on the GPR surrogate energy surface. Also, several of the systems in the dataset involve two separate molecular fragments and in a few cases, the saddle point calculation moves them away from each other until the calculations are aborted. The reason for this is that the direction of the minimum mode then corresponds to the uniform displacement of the two fragments with little change in their internal structure. The energy decreases as the fragments move closer together and so the saddle point search moves the fragments in the opposite direction further apart so as to climb uphill on the energy surface. Similar behaviour is observed in the Sella calculations of these systems. However, in most cases where this occurs, the sideways minimization of the energy eventually leads to the search path turning around and the two fragments then eventually come together in a chemical reaction. However, in a few cases, this does not happen and the GPR-dimer calculation is eventually aborted after a maximum number of iterations has been reached. When the initial configuration of the atoms consists of separated molecular fragments, it could make more sense to start with a local minimization of the energy to bring the fragments to an optimal (relative) position, and then start the saddle point search from a configuration of the atoms that is only slightly perturbed from the minimum energy configuration. Here, we have chosen to work with the input as published by [23].
@@
The application of a saddle point search method that is designed to move uphill in energy is actually questionable when the initial state consists of two molecular fragments since the uphill climb in energy will likely move the two fragments apart. An NEB calculation of the minimum energy path for such a case does not run into this kind of a problem, but then the final state of the reaction needs to be specified. Furthermore, the energy surface of the two fragments may be poorly described with a harmonic approximation. Then, harmonic transition state theory is not expected to provide a good estimate of the rate. In principle, a free energy calculation should be applied in such cases, and, an optimal dividing surface should be determined by maximizing the free energy [26].
@@
An important aspect of the use of a surrogate energy surface is not to venture too far from the region of the data points in the training set since extrapolation is not reliable. This is implemented by early stopping criteria in the GPR-dimer saddle point searches [31, 30]. . The number of times early stopping criteria are applied correlates with longer searches, as the structures added to the training data due to the early stopping are not typically near a saddle point. Here, Euclidean distances have been used to decide on early stopping, but we expect that a Hausdorff distance measure could work better in tandem, and this will be tested in future refinements of the method. Additionally, the Euclidean distance measure is not sensitive to changes across chemical species, which can lead to chemically unfeasible structures being marked as within the cutoff.
@@
The GPR-dimer saddle point search on the surrogate surface can lead to configurations of the atoms that have high energy. The NWChem software used for the HF calculations performs checks on the input structures before calculating the energy and forces, and these are triggered even when the GPR is not used. While the GPR-dimer explores unphysical structures, high energy and atomic forces are in principle not problematic as the GPR can successfully handle such regions. However, NWChem often terminates when atoms come too close and this prevents the GPR-dimer from navigating these regions. We hypothesise that by providing some energy/force value, even a large one representing extreme repulsion, as a pseudo-output from the electronic structure calculation through a wrapper in ASE or EON would allow the GPR-dimer to successfully converge to a saddle point in such cases, while here they represent a failure.
@@
Another improvement that is being pursued for future refinements of the method is the pruning of input data, namely, the elimination of some of the configurations that have been calculated with the electronic structure method early in the search and are far from the saddle point. This will help reduce the computational cost of the matrix inversion and the memory requirement.
@@
In future work we intend to apply these and other methods to larger and more diverse benchmark sets that are not biased towards any particular method. The results of such studies will be reported in an ongoing collaboration to extend the OptBench suite of benchmarks for finding saddle points and minimum energy paths of thermally activated transitions [5].
@@ 5 Conclusions
The results presented here demonstrate that saddle point searches for molecular reactions can be performed efficiently using Cartesian coordinates despite the large disparity in the stiffness of the molecular degrees of freedom if the calculation makes use of GPR to generate a surrogate energy surface. The MMF method is used here in combination with a dimer estimate of the minimum mode. The use of a GPR reduces the number of electronic structure calculations needed to reach convergence by an order of magnitude.
The calculations are carried out for a dataset of 500 systems presented previously by [23]. It consists of guesses for saddle points and by construction only includes cases where calculations using Sella are successful in that some saddle point is obtained. A comparison is made between the performance of the GPR-dimer method using Cartesian coordinates and the algorithm in Sella based on internal coordinates measured in terms of the number of HF calculations required for convergence.
The performance of the two methods is found to be similar. However, in several cases the Sella calculation bypasses the saddle point that is closest to the initial guess and converges instead on some rearrangement of the atoms in the final state, such as a rotation of a methyl group. This demonstrates the importance of checking the saddle point configuration of atoms obtained in a single endpoint calculation. This can be done efficiently by using the NEB method to find the minimum energy path as it can reveal whether the saddle point found is adjacent to the initial state configuration or whether another saddle point is present in between, as is illustrated in Figure 4.
While the use of a GPR presents a significant overhead for the saddle point search, the efficient implementation presented here in C++ leads to a reduction in the overall computational time even though the electronic structure calculations are carried out here at a low HF level. When higher level electronic structure calculations are used, the overhead will be an even smaller fraction of the total computational effort. The computational effort in the generation of the GPR surrogate energy surface scales rapidly with the dimensionality of the system and the number of electronic structure results included in the training set. Further optimisation is required in order to be able to apply this approach to systems with a large number of movable atoms. Pruning of the input data to drop points far from the saddle point region is one option and the use of only a subset of the atom coordinates of a large system focusing thereby on the region of interest is another, possibly followed by later refinement of the saddle point found with respect to all coordinates.
6 Acknowledgements
@@
The authors confirm that the data supporting the findings of this study are available within the article and/or its supplementary materials. The supplementary material includes: a single repository, which is tagged. These include pinned versions of ASE, EON (from https://theochemui.github.io/eOn/), the GPR dimer code, and a copy of the spack recipes [11] repository for building NWChem used for the HartreeFock calculations along with spinout utilities from Wailord [13]. A conda ecosystem based environment managed by pixi is provided to further facilitate reproducibility, with the modules either pinned to versions or bundled within the same repository via gitsubrepo following best practices for literate programming [15, 16].
Data related to the results presented in this article and instructions on the generation thereof are available in [14] hosted on the Materials Cloud Archive [48].